# Time-Transformer: Integrating Local and Global Features for Better Time Series Generation

### 18 Dec 2023 v1

@@ 5.3.4 Model Application: Data Augmentation
Time-Trans. 0.9802 0.8178 0.9163 0.9945 0.9345
We do not list TimeGAN because training data here are insufficient to train this model. The results on the other three datasets can be found in D.4. One crucial problem of imbalanced classification is low recall rates, caused by under-represented positive class in the training set. This can result in negative prediction dominance. Repeating the minority class can improve the recall at the cost of precision, which reflects the aforementioned overfitting problem. However, augmenting the training set with synthetic data from our model shows better results with respect to all metrics suggesting that it can aid in solving this kind of problem.
Using three balanced UCR datasets ($<$ 2000 time series), we investigate how 3 generative models can improve the deep learning training for small datasets. We extend these datasets with 25%, 50%, 75%, and 100% of their size and use them to train the same MLP. Figure 6 shows the change of accuracy and area under ROC and PRC curves. The classifier performs better with more synthetic training data, and best with those from our model. Detailed results are provided in D.5.
Figure 6: Classification results on augmented datasets with: Time-Trans.(green), RTSGAN(blue), and TimeVAE(orange). Higher results are better.
5.3.5 Ablation Study & Comparison with Sequential Design
Using the ‘Music’ dataset, we investigate how each component of the model contributes to the final result. We (1) use only a De-convolutional block, (2) add only a TCN block after (1), (3) add only a Transformer block after (1), and (4) add our proposed Time-Trans. after (1). Additionally, to test our claim that parallel structure is better than sequential combination, we run an extra experiment where we add sequential combination of TCN and Transformer (Trans.) after (1). Results are shown in Table 5(results of other two local-global datasets are in D.7). From the FID column we see a clear improvement when adding either a TCN or a Transformer. The sequential combination also improves the performance. Finally, our proposed parallel structure Time-Trans. achieves an approximate 40% improvement compared to the sequential one. The other two columns also reflect similar situations. These results indicate the strong ability of our proposed model to generate time series.
@@ 5.3.5 Ablation Study & Comparison with Sequential Design
Trans. 0.815±0.039 0.479±0.009 0.027±0.006
TCN+Trans. (Seq.) 0.627±0.031 0.455±0.021 0.022±0.008
Time-Trans. 0.395±0.013 0.160±0.054 0.017±0.009
@@ 6 Limitations & Border Impacts
Although we showed the strength of our model from multiple perspectives, we acknowledge its limitations and propose some corresponding future works. First, as our model handles unsupervised generation, changes would need to be made for a conditional generation task, to produce user-defined synthetic time series data. Second, aiming at data augmentation, our model requires access to complete time series data. Expending it to work with incomplete data and address time series imputation challenges would offer significant benefits. Apart from the positive impacts showed in 5.3.4, our model does have possible negative social impact if used unethically. E.g. realistic fake time series data might be produced for fake news to mislead people. It is important to follow ethical practices when using the model.
@@ 7 Conclusion
We introduced a novel time series generative model (Time-Transformer AAE), which contains an adversarial autoencoder (AAE) and a key component (Time-Transformer) in the decoder. Via a layer-wise parallelization and a bidirectional cross attention, Time-Transformer AAE exploits the learning ability of Temporal Convolutional Networks and Transformer in extracting local features and building global interaction respectively. We experimentally showed its effectiveness in generating time series data. Additionally, by training on different sizes of data, we show the proposed model can achieve competitive performance to SOTA methods, even with less training data. Furthermore, we showed the model’s utility in imbalanced classification and small dataset augmentation problems.
Acknowledgements: This work is partially supported by NHMRC Development Grant Project 2000173.

### 28 Jan 2024 v3

@@ 5.3.4 Model Application - Data Augmentation:
Time-Trans 0.9802 0.8178 0.9163 0.9945 0.9345
We do not list TimeGAN because training data here are insufficient to train this model. The results on the other three datasets can be found in [38]. One crucial problem of imbalanced classification is low recall rates, caused by an under-represented positive class in the training set. This can result in negative prediction dominance. Repeating the minority class can improve the recall at the cost of precision, which reflects the aforementioned overfitting problem. However, augmenting the training set with synthetic data from our model shows better results with respect to all metrics suggesting that it can aid in solving this kind of problem.
Using the balanced datasets ($<$ 2000 time series), we investigate how 3 generative models can improve the deep learning training for small datasets. We extend these datasets with 25%, 50%, 75%, and 100% of their size and use them to train the same MLP. Figure 6 shows the change of accuracy and area under ROC and PRC curves. The classifier performs better with more synthetic training data, and best with those from our model. Detailed results are provided in [38].
Figure 6: Classification results on augmented datasets with: Time-Trans(green), RTSGAN(blue), and TimeVAE(orange). Higher results are better.
5.3.5 Ablation Study & Comparison with Sequential Design:
Using the ‘Music’ dataset, we investigate how each component of the model contributes to the final result. We (1) use only a De-convolutional block, (2) add only a TCN block after (1), (3) add only a Transformer block after (1), and (4) add our proposed Time-Trans after (1). Additionally, to test our claim that a parallel structure is better than a sequential combination, we run an extra experiment where we add a sequential combination of TCN and Transformer (Trans) after (1). Results are shown in Table 5(results of the other two local-global datasets are in [38]). From the FID column we see a clear improvement when adding either a TCN or a Transformer. The sequential combination also improves the performance. Finally, our proposed parallel structure Time-Trans achieves an approximate 40% improvement compared to the sequential one. The other two columns also reflect similar situations. These results indicate the strong ability of our proposed model to generate time series.
@@ 5.3.5 Ablation Study & Comparison with Sequential Design:
Trans 0.815±0.039 0.479±0.009 0.027±0.006
TCN+Trans (Seq.) 0.627±0.031 0.455±0.021 0.022±0.008
Time-Trans 0.395±0.013 0.160±0.054 0.017±0.009
@@ 6 Limitations & Border Impacts
Although we showed the strength of our model from multiple perspectives, we acknowledge there are some limitations and propose future work to mitigate these. First, as our model handles unsupervised generation, changes would need to be made for a conditional generation task, to produce user-defined synthetic time series data. Second, aiming at data augmentation, our model requires access to complete time series data. Expanding it to work with incomplete data and address time series imputation challenges would offer significant benefits. Apart from the positive impacts showed in 5.3.4, our model may have possible negative social impact if used unethically (E.g. realistic fake time series data in news may be used to mislead people). It is important to follow ethical practices when using the model.
@@ 7 Conclusion
We introduced a novel time series generative model (Time-Transformer AAE), which contains an adversarial autoencoder (AAE) and a key component (Time-Transformer) in the decoder. Via a layer-wise parallelization and a bidirectional cross attention, Time-Transformer AAE exploits the learning ability of Temporal Convolutional Networks and Transformer in extracting local features and building global interaction respectively. We experimentally showed its effectiveness in generating time series data. Additionally, by training on different sizes of data, we showed the proposed model can achieve competitive performance to SOTA methods, even with less training data. Furthermore, we showed the model’s utility in imbalanced classification and small dataset augmentation problems.
Acknowledgements: This work is partially supported by NHMRC Development Grant Project 2000173 and The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.